Assume that the real exchange rate is equal 4(US goods/CAN goods). Also, assume that P
US = 1 and P
C AN = 2
as well as that there are no transportation costs and all goods are tradable. Suppose that there is a
Canadian investor looking to invest C$1, 000 by buying goods in one market and selling them in
the other. What is the maximum profit the investor can make if she takes advantage of the arbitrage
opportunity (express profits in C$)?
1
Expert's answer
2019-12-13T10:20:01-0500

Macroeconomics


We need to find the maximum profit.


Solution:


Given,


Real exchange rate = 4


Nominal exchange rate =

"E =R \times \frac {P_{US}}{P_{CAN}}"

"E = 4 \times \frac {1}{2} = 2"

An individual amount invested by Canadian investor = C$1, 000


This investment is equal to $2000


The price of the good in US is PUS = $1


So, the numbers of good that can be purchased is "=\frac {\$2000}{\$1} = 2000 \space units"


if these goods will be sold in Canada then the individual will get "C$2*2,000 = C$4,000”


So, here the maximum profit is = "4,000 \u2013 1,000= 3000 >0"